OpenStudy (anonymous):

Draw the angle given in degrees on the unit circle where 0 radians corresponds to the positive portion of the horizontal axis: -16pi/7

OpenStudy (anonymous):

anyone ?

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

plz help me

OpenStudy (jdoe0001):

\(\bf \cfrac{16}{7}\implies \cfrac{7+7+2}{7}\implies \cfrac{\cancel{7}}{\cancel{7}}+\cfrac{\cancel{7}}{\cancel{7}}+\cfrac{2}{7}\implies 1+1+\cfrac{2}{7} \\ \quad \\ 2+\cfrac{2}{7} \\ \quad \\ thus\qquad -\cfrac{16}{7}\implies -2\pi -\cfrac{2\pi }{7}\) notice, a full CLOCKWISE cycle is \(-2\pi\) so just draw the remaining after the cycle and then get the CLOCKWISE angle

OpenStudy (jdoe0001):

shoot... actually that's backwards

OpenStudy (jdoe0001):

notice, a full CLOCKWISE cycle is −2π so just draw the remaining after the cycle and then get the COUNTER-CLOCKWISE angle

OpenStudy (jdoe0001):

anyhow \(\bf -\cfrac{16\pi }{7}\implies -2\pi -\cfrac{2\pi }{7}\)